Grade 12Physical Chemistry

when 6 gram of non volatile solute is added to 60 gram of benzene it elevates the boiling point by 1 degree celcius if the value ko Kb for benzene is 2.53K/kg then what is the molar mass of non volatile solut

aayushi
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
To solve we use formula,  ΔTb = Kb × i × m
Where ΔTb is change in boiling temperature = 1°C or 1K
i  is van’t Hoff factor =1  (solute does not dissociate )
Kb = molal boiling point elevation constant = 2.52 K kg/mol
m is molality of solute or  m= moles of solute/mass of solvent in kg 
 
we need molality in terms of molar mass

 m= w2M2×w1
w2 = mass of solute = 6g
M2 = molar mass of solute
w1 = mass of solute in kg = 60 g or 0.06kg
 
Substituting in equation 1 and rearranging
 We get answer as 253 gm/mol
M2 = i×Kb×w2ΔTb×w1

M2 = 1×2.52×101×0.1M2= 252g/mol